OpenStudy (anonymous):

6.Given the quadratic equation mx^2-2mx-1=6x-m has two different roots,find the range of values of m. @ganeshie8 @SolomonZelman @UnkleRhaukus

OpenStudy (solomonzelman):

I am not very sure how to do this, I would start of doing this.... (completing the square) \(\large\color{slate}{ mx^2-2mx-1=6x-m }\) \(\large\color{slate}{ mx^2-2mx-6x=1 -m }\) \(\large\color{slate}{ mx^2-x(2m+6)=1 -m }\) \(\large\color{slate}{ m(x^2-x\frac{2m+6}{m})=1 -m }\) \(\large\color{slate}{ x^2-x\frac{2m+6}{m}=\frac{1 -m }{m}}\) \(\large\color{slate}{ x^2-x\frac{2m+6}{m}+\frac{(m+3)^2}{m^2}=\frac{1 -m }{m}+\frac{(m+3)^2}{m^2}}\) \(\large\color{slate}{ (x-\frac{m+3}{m})^2=\frac{1 -m }{m}+\frac{(m+3)^2}{m^2}}\) \(\large\color{slate}{ (x-\frac{m+3}{m})^2=\frac{(1-m)^2+(m+3)^2}{m^2}}\) \(\large\color{slate}{ (x-\frac{m+3}{m})^2=\frac{1-2m+m^2+m^2+6m+9}{m^2}}\) \(\large\color{slate}{ (x-\frac{m+3}{m})^2=\frac{2m^2+4m+10}{m^2}}\)

OpenStudy (anonymous):

answer is \[m>-\frac{ 9 }{ 7 }\]

OpenStudy (solomonzelman):

why?

OpenStudy (solomonzelman):

(not necessarily saying it is wrong, just want to see what you have done to obtain it)

OpenStudy (anonymous):

no,i'm telling u the given answer.

OpenStudy (solomonzelman):

Oh, idk how they did it.

OpenStudy (solomonzelman):

(this is my gap in math... I would really like to see how someone else does this prob so that I know for future reference)

OpenStudy (solomonzelman):

sorry that I couldn't help-:(

OpenStudy (eric_d):

use b^2-4ac>0 @MARC_

OpenStudy (eric_d):

before that

OpenStudy (eric_d):

rearrange it in this form ax^2+bx+c>0

OpenStudy (unklerhaukus):

A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] The radicand of the quadratic formula is called the Discriminant \(\Delta =b^2-4ac\) if \(\Delta=0\) there is only one solution to the quadratic equation if \(\Delta>0\) there are two solutions if \(\Delta<0\) there are no real solutions So, rearrange you equation of into the form \[ax^2+bx+c=0\] For two real roots the discriminate of this must greater then zero

OpenStudy (anonymous):

\[mx^2-2mx-6x+m-1=0\]

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

read what unklerhakus wrote

OpenStudy (dan815):

mx^2-2mx-1=6x-m mx^2+(-2m-6)x-1+m=0 so under the sqrtb^2-4ac you have b^2-4ac=(-2m-6)^2-4*m*(-1+m) b^2-4ac>0 so that its not repeating or imaginary roots (-2m-6)^2-4*m*(-1+m) > 0 solve for m

OpenStudy (anonymous):

8m^2+28m+60>0

OpenStudy (anonymous):

Is this correct so far ? @dan815

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

28m+36>0

OpenStudy (anonymous):

m>-9/7

OpenStudy (anonymous):

Got it finally !

ganeshie8 (ganeshie8):

looks good!

OpenStudy (anonymous):

Thank you @dan815 @UnkleRhaukus @ganeshie8 @eric_d @SolomonZelman